john, twelve years old, is three times as old as his brother. how old will john be when when he is twice as old as his brother?
Answer by jorel1380(3719)

If John is 12,1/3rd of his age is 4, which is his brother's age.
12+x=2(4+x)
12+x=8+2x
4=x
In 4 more years, John will be twice his brother's age.